A cylinder has a radius of 5 in. If the volme of the cylinder is 250π in.³, what is the height of the cylinder?
[Hint: The volume of a cylinder is given by V = πr²h.]

Answer by ankor@dixie-net.com(12701)

A cylinder has a radius of 5 in. If the volume of the cylinder is 250π in.³, what is the height of the cylinder?
:
[Hint: The volume of a cylinder is given by V = πr²h.]
:
Substituting in the given formula:
pi*5^2*h = 250
:
pi*25*h = 250; squared 5
:
78.5398h = 250; multiplied by pi
:
h = 250/78.5398; divided both sides by 78.5398
:
h = 3.1831 in is the height
:
:
Check solution with h = 3.1831 using a calc:
V = pi*5^2*3.1831
V = 250.00
:
That wasn't so hard, right?